Let demand be given by Q = 150 - P + 2Y. This is the same for all problems of this type. Let r = 10%. Let Y = 50 across both periods. Let MC = 30. Let reserves = 200. Consider the basic two-period model. What is consumption of the resource in the present?

94.29

105.71

107.14

111.11

none of the above
Answer #1

Answer option 2)

Q= 150-P+2*50 = 250-P

P= 250-Q

MNB = Marginal net benefit

= P-MC

at eqm,

MNB1= (1/1+r)*MNB2

250-30-Q1 = (250-30-Q2)/(1.1)

220-Q1 = (220-Q2)/1.1

Q1+ Q2= 200

solving two equations

Q1*= 105.714